OpenStudy (anonymous):

SHOW THE EQUATION OF A LINE, IN SLOPE-INTERCEPT FORM, THAT PASSES THROUGH THE POINT (-3,5) AND HAS A SLOPE OF -4

OpenStudy (cybershadow):

well do u know what the slope intercept form of the line is?

OpenStudy (anonymous):

I KNOW THE EQUATION IS y = mx+b.

OpenStudy (anonymous):

your caps lock hurts my eyes, pls relieve me of my torture

OpenStudy (cybershadow):

Yeap and where m is the slope right? also b is the y -intercept now as the line passes through (-3,5) , it must satisfy the y = mx + c equation put the values of the m , x, and y (these are given) and find the y-intercept? can u do it?

OpenStudy (cybershadow):

O_O i never knew i could be this patient .-..

OpenStudy (cybershadow):

so much for a "Yes i can"/"No i have this doubt" Answer o_O

OpenStudy (anonymous):

y=-4x-7

OpenStudy (cybershadow):

well thats correct !

OpenStudy (cybershadow):

Just for the googlers.... here is the slope form of the line y = mx + c (where , x and y are x,y coordinates respectively and c is the y-intercept) now we have been given m = -4 y = 5 and x = -3 just put it in above equation that will make it 5 = -4(-3) + c c = 5-12 = - 7 therefore equation becomes , y = -4x-7